Circular No. 6890

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bp IN NGC 6495 G. M. Hurst, Basingstoke, England, reports the discovery by Mark

Armstrong, Rolvenden, of an apparent supernova (mag 14.8) on unfiltered CCD images taken on Apr. 29.074, 29.101 and 29.128 UT with a 0.26-m Schmidt Cassegrain telescope in the course of the U.K. Nova/Supernova Patrol. The new star was reported at α = 17^{h}54^{m}50^{s}.71, δ = +18^{o}19′49′′.3 (equinox 2000.0), or ∼ 15′′ due north of the nucleus of NGC 6495. No star was recorded at this position on a CCD frame taken by Armstrong on Mar. 20.283 (limiting mag 16.7). Hurst estimated mag 14.7 from a CCD frame obtained by T. Boles, Wellingborough, on Apr. 29.135 in strong twilight. S. Nakano, Sumoto, Japan, forwards the following position end figures measured by A. Sugie at Dynic Astronomical Observatory, from CCD frames taken under poor conditions on Apr. 30.54 (V = 14.9–15.1): 50^{s}.72, 50′′.2, which is 1′′ west and 13′′ north of the galaxy’s center.

F. Patat, European Southern Observatory (ESO); and M. Maia, Observatorio Nacional, Rio de Janeiro, report on their observations of SN 1998bp with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph; 10-min exposure, range 360–750 nm, resolution 0.5 nm) obtained on Apr. 30.38 UT at La Silla: “A full reduction of the CCD spectrogram shows that this object is indeed a supernova. The spectral features are consistent with a peculiar type-Ia object, roughly around maximum light. The usual Si II feature (635.5 nm) is much less deep than in normal type-Ia supernovae, if compared with all of the other spectral lines. The spectrum is dominated by P-Cyg profiles, superimposed on a slightly blue continuum. The most prominent features are identified as Ca II H and K, Si II, Mg II, and S II. From the spectrophotometry we estimate B −V ∼ +0.4 and V −R ∼ +0.4, which are rather red for a normal type-Ia supernova at this phase. Since the host galaxy is elliptical and there is no evident signature of Na I D absorption in the spectrum, the red observed color is intrinsic. The expansion velocities measured from the minima of Ca II H and K, Mg II 448.1-nm, and Si II 635.5-nm are 13 600, 13 700, and 12 800 km/s, respectively. The recession velocity of the parent galaxy, deduced from the emission peak of the Si II 635.5-nm line, is 4500 km/s.”

COMET C/1998 H1 (STONEHOUSE) Visual m_{1} estimates: Apr. 27.83 UT, 10.4 (K. Hornoch, Lelekovice,

Czech Rep., 0.35-m refl.); 28.92, 10.4 (M. Reszelski, Szamotuly, Poland, 20×60 binoc.); 29.95, 10.4 (A. Baransky, Kiev, Ukraine, 0.20-m reflector).

1998 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green